# Robust Point Matching with Distance Profiles

### v2

@@ Abstract
While matching procedures based on pairwise distances are conceptually appealing and thus favored in practice, theoretical guarantees for such procedures are rarely found in the literature. We propose and analyze matching procedures based on distance profiles that are easily implementable in practice, showing these procedures are robust to outliers and noise. We demonstrate the performance of the proposed method using a real data example and provide simulation studies to complement the theoretical findings.
@@ 1 Introduction
Matching is one of the most fundamental tasks of modern data science, which can be found in numerous applications. For instance, the goal of shape correspondence problems [Van Kaick et al., 2011] in computer graphics is to match corresponding parts of different geometric shapes, often obtained by 3D scanners or imaging equipment used for Magnetic Resonance Imaging (MRI) or Computed Tomography (CT). Graph matching aims to find corresponding nodes between two or more graphs representing structured data, such as neuronal connectivity in human brains [Sporns et al., 2005] or social networks [Borgatti et al., 2009].
@@
A formal mathematical framework for matching [Mémoli, 2011] conceptualizes objects as metric measure spaces, which are essentially probability measures defined on suitable metric spaces. For instance, a smooth 3D surface can be viewed as a uniform measure supported on the surface in the Euclidean space $\mathbb{R}^{3}$. In practice, of course, objects are not fully available and thus approximated by suitable point clouds, say, $\{X_{1},\ldots,X_{n}\}$ and $\{Y_{1},\ldots,Y_{m}\}$ on $\mathbb{R}^{d}$, respectively, and the practical goal is to match these point clouds. The most standard approach is quadratic matching which formulates and solves a quadratic program based on pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$. When $n=m$, quadratic matching finds a one-to-one correspondence that aligns the pairwise distances as closely as possible, which is formulated as a Quadratic Assignment Problem (QAP) [Koopmans and Beckmann, 1957, Loiola et al., 2007]. When $n\neq m$, the notion of correspondence is relaxed to soft assignments or couplings, leading to the formulation of the Gromov-Wasserstein (GW) distance [Mémoli, 2011], which has been increasingly used in applications, such as word alignment [Alvarez-Melis and Jaakkola, 2018], statistical inference [Brécheteau, 2019, Weitkamp et al., 2022], and generative modeling [Bunne et al., 2019, Hur et al., 2024]. On the practical side, however, quadratic matching requires solving a quadratic program that is NP-hard in general, and thus needs to rely on several methods, such as convex relaxation [Aflalo et al., 2015], semidefinite relaxation [Zhao et al., 1998, Kezurer et al., 2015], spectral relaxation [Leordeanu and Hebert, 2005], or entropic regularization [Rangarajan et al., 1999, Solomon et al., 2016, Peyré et al., 2016].
@@
Matching procedures based on the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are invariant to rigid transformations, which is desirable in practice. For instance, in imaging problems, the images of the objects of interest are often obtained from different angles and thus need to be matched up to rigid transformations. However, such procedures are difficult to analyze theoretically and the literature has been mostly establishing theoretical guarantees for other matching procedures using $\|X_{i}-Y_{j}\|_{2}$’s, which aim to match the points based on the proximity. [Collier and Dalalyan, 2016, Kunisky and Niles-Weed, 2022, Wang et al., 2022] establish theoretical results on recovering the ground truth matching for the procedures based on $\|X_{i}-Y_{j}\|_{2}$’s under some noise and [Galstyan et al., 2022, Minasyan et al., 2023, Galstyan and Minasyan, 2023] show matching guarantees in the presence of outliers. Though Theorem 2 of [Wang et al., 2022] is concerned with a procedure using the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, it requires solving a nontrivial minimization over permutations, which is computationally challenging, and the result does not consider the presence of outliers.
@@
Motivated by the above challenges, this paper theoretically establishes outlier robustness and noise stability for a matching procedure based on the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, which is easily implementable in practice. The proposed procedure associates each point with the set of distances from it to other points, namely, $X_{i}$ and $Y_{j}$ are associated with the sets of distances to the other points $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$, which we call the distance profiles of $X_{i}$ and $Y_{j}$, respectively. The matching procedure is to find for each $X_{i}$, the point $Y_{j}$ having the most similar distance profile to that of $X_{i}$, where the similarity is measured by a suitable quantity between the sets $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$. The idea of associating a point with suitable invariants based on the distances to other points from it has a long history in the object matching literature, which includes shape contexts [Belongie et al., 2002, Ruggeri and Saupe, 2008], integral invariants [Manay et al., 2006], invariant histograms [Brinkman and Olver, 2012], and the local distribution of distances [Mémoli, 2011]. However, theoretical results on robustness of such methods are rarely found in the literature. In the context of random graph matching, [Ding et al., 2021, Ding et al., 2023] propose and analyze a degree profile matching procedure based on the adjacency matrices $(A_{i\ell})_{i,\ell=1}^{n}$ and $(B_{j\ell})_{j,\ell=1}^{n}$ of certain correlated random graphs. However, since their analysis focuses on random graph models, such as Erdös-Rényi and correlated Wigner models, where the edges are independently generated, $(A_{i\ell})_{i,\ell=1}^{n}$ are mutually independent, which is substantially different from our setting where the distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ are not mutually independent.
@@
We show the outlier robustness and noise stability of the matching procedure based on distance profiles, which is easy to implement in practice. Though the main principle of the proposed procedure, namely, matching the points via certain distance-based invariants assigned to them, has already been proposed and rediscovered multiple times in the literature as we mentioned earlier, there has been no theoretical analysis on the outlier robustness and noise stability of this procedure to the best of our knowledge. Under certain probabilistic settings, we prove that the proposed matching procedure is successful with high probability even in the presence of outliers and noise. We demonstrate the performance of the proposed method using a real data example and provide simulation studies to complement the theoretical findings.
@@
For any $k\in\mathbb{N}$, let $[k]$ denote $\{1,\ldots,k\}$, $\Delta_{k}=\{(p_{1},\ldots,p_{k})\in\mathbb{R}_{+}^{k}:\sum_{i=1}^{k}p_{i}=1\}$ denote the probability simplex, and $\mathcal{S}_{k}$ denote the set of all permutations on $[k]$. For $a,b\in\mathbb{R}$, let $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. Let $\mathscr{P}(\mathbb{R}^{d})$ denote the set of all Borel probability measures defined on $\mathbb{R}^{d}$ and for any $z\in\mathbb{R}^{d}$, let $\delta_{z}\in\mathscr{P}(\mathbb{R}^{d})$ denote the Dirac measure at $z$. Lastly, we denote by $W_{1}$ the Wasserstein-1 distance.
@@ 1.1 Distance Profile Matching
Algorithm 1 describes the distance profile matching procedure that we propose to study. First, notice that Algorithm 1 is implementable without knowing the points $X_{i}$’s and $Y_{j}$’s as long as the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are provided. The discrepancy measure $\mathcal{D}$ quantifies the similarity between the distance profiles. Therefore, for each $X_{i}$, the procedure (1.2) is finding $Y_{j}$ whose distance profile is closest to that of $X_{i}$ under $\mathcal{D}$.
@@
The output $\pi$ of Algorithm 1 represents the matching: $X_{i}$ is matched to $Y_{\pi(i)}$. Note that $\pi$ may not be injective or surjective, meaning that some points may not be matched or some points may be matched to multiple points. [^1] If $n=m$ and one wants to find a one-to-one correspondence, Algorithm 1 can be modified to output a permutation as we discuss in Section 3. The motivation behind this is simple: (1) in the presence of outliers, some points can be ignored or matched arbitrarily, and (2) similar points can be matched to the same point. The underlying objects that the points $X_{i}$’s and $Y_{j}$’s approximate, say, probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{d}$, respectively, are not identical in practice. Usually, they coincide up to several outlier parts, which we can formally state as $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$ for some $\lambda\in(0,1)$ and $\rho,\mu_{o},\nu_{o}\in\mathscr{P}(\mathbb{R}^{d})$. Here, $\lambda\rho$ represents the common part of $\mu$ and $\nu$ that we want to match, while $(1-\lambda)\mu_{o}$ and $(1-\lambda)\nu_{o}$ are the outlier parts of $\mu$ and $\nu$, respectively. We are mainly interested in matching $X_{i}$’s and $Y_{j}$’s from the common part, while ignoring or arbitrarily matching the points from the outlier parts.
@@
The threshold $\rho$ is an optional parameter that can be used to obtain the subset $I\subset[n]$ containing the indices $i$’s such that the discrepancy $D(i,\pi(i))$ between the distance profiles of $X_{i}$ and its match $Y_{\pi(i)}$ is below $\rho$. Providing a small threshold leads to a more conservative matching that only focuses on the points that are closely matched, which can be used for other downstream tasks. For instance, one can solve the orthogonal Procrustes problem using the matched pairs $\{(X_{i},Y_{\pi(i)})\}_{i\in I}$ to find the optimal rotation and translation aligning them as closely as possible, which theoretically works with just $|I|=d$ pairs in $\mathbb{R}^{d}$ but practically requires more pairs, say, $|I|=O(d)$, due to potential noise or degeneracy. The obtained transformation can then be used as an initialization of the registration procedures, such as the Iterative Closest Point (ICP) algorithm [Besl and McKay, 1992], to find the optimal transformation that aligns the whole point clouds.
@@
The Wasserstein-1 distance between two sets of points in $\mathbb{R}$ admits a closed-form expression based on the corresponding cumulative distribution functions, which can be computed efficiently [Peyré and Cuturi, 2019].
@@
1.2 Motivating Application
Before presenting the main theoretical results, we demonstrate how the proposed distance profile matching procedure performs in practice using a real data example. We apply the distance profile matching method to the shape correspondence using CAPOD data [Papadakis, 2014]. Figure 1(a) and Figure 1(d) show the two 3D images representing dolphins with different poses, from which we obtain point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ as shown in Figure 1(b) and Figure 1(e), respectively, where $n=m=285$ in this example. We want to match the points corresponding to the same part of the dolphins, rather than finding a one-to-one correspondence. As explained earlier, it is reasonable to match similar points to the same point. For instance, there are many points concentrated around the nose—called the rostrum—of the source dolphin, which are expected to be matched to any points concentrated around the rostrum of the target dolphin. Hence, we first divide the dolphins into several components. Here, we use $k$-means clustering with $k=7$, which indeed cluster points based on the body parts. Figure 1(e) shows the target points $Y_{j}$’s colored by $s(j)$’s, where $s(j)\in\{1,\ldots,7\}$ is the index of the component that $Y_{j}$ belongs to. Now, we run Algorithm 1 with $\mathcal{D}=W_{1}$ and obtain the output $\pi\colon[n]\to[m]$. For each $X_{i}$, we verify if the part that it belongs to coincides with the part that its match $Y_{\pi(i)}$ belongs to, which is indexed by $s(\pi(i))$. To this end, Figure 1(b) shows the source points $X_{i}$’s colored by $s(\pi(i))$’s. We can observe that the proposed procedure recovers the corresponding regions mostly correctly. For instance, for $X_{i}$’s corresponding to the rostrum of the source dolphin, their colors $s(\pi(i))$’s (green) indicate that they are matched to the points corresponding to the rostrum of the target dolphin. Similarly, the points corresponding to the head (lime), tail (red), dorsal fin (blue), and peduncle (magenta) are mostly matched correctly, while there are some mismatches, for instance, among the points around the pectoral fin.
Figure 1(c) and Figure 1(f) repeat the same procedure after removing 16 points around the fin of the target dolphin in Figure 1(e) and clustering again with $k=6$. Figure 1(c) shows that the matching result is similar to the previous one. In this case, though the points around the fin of the source dolphin do not have corresponding points in the target dolphin and thus we may ignore matching them, the proposed procedure still matches them to the points around the peduncle (magenta) of the target dolphin, which is a cluster adjacent to the removed fin.

### v5

@@ Abstract
We show the outlier robustness and noise stability of practical matching procedures based on distance profiles. Although the idea of matching points based on invariants like distance profiles has a long history in the literature, there has been little understanding of the theoretical properties of such procedures, especially in the presence of outliers and noise. We provide a theoretical analysis showing that under certain probabilistic settings, the proposed matching procedure is successful with high probability even in the presence of outliers and noise. We demonstrate the performance of the proposed method using a real data example and provide simulation studies to complement the theoretical findings. Lastly, we extend the concept of distance profiles to the abstract setting and connect the proposed matching procedure to the Gromov-Wasserstein distance and its lower bound, with a new sample complexity result derived based on the properties of distance profiles. This paper contributes to the literature by providing theoretical underpinnings of the matching procedures based on invariants like distance profiles, which have been widely used in practice but have rarely been analyzed theoretically.
@@ 1 Introduction
Matching is one of the most fundamental tasks of modern data science, which can be found in numerous applications. For instance, the goal of shape correspondence problems [37] in computer graphics is to match corresponding parts of different geometric shapes, often obtained by 3D scanners or imaging equipment used for Magnetic Resonance Imaging (MRI) or Computed Tomography (CT). Graph matching aims to find corresponding nodes between two or more graphs representing structured data, such as neuronal connectivity in human brains [35] or social networks [6].
@@
A formal mathematical framework for matching [26] conceptualizes objects as metric measure spaces, which are essentially probability measures defined on suitable metric spaces. In practice, of course, objects are not fully available and thus approximated by suitable point clouds, say, $\{X_{1},\ldots,X_{n}\}$ and $\{Y_{1},\ldots,Y_{m}\}$ on $\mathbb{R}^{d}$, respectively, and the practical goal is to match these point clouds. Quadratic matching is one of the most standard approaches to this problem, which formulates and solves a quadratic program based on pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$. When $n=m$, quadratic matching finds a one-to-one correspondence that aligns the pairwise distances as closely as possible, which is formulated as a Quadratic Assignment Problem (QAP) [21, 24]. When $n\neq m$, the notion of correspondence is relaxed to soft assignments or couplings, leading to the formulation of the Gromov-Wasserstein distance [26], which has been increasingly used in applications, such as word alignment [2], statistical inference [7, 41], and generative modeling [9, 19]. On the practical side, however, quadratic matching requires solving a quadratic program that is NP-hard in general, and thus needs to rely on several methods, such as convex relaxation [1], semidefinite relaxation [42, 20], spectral relaxation [23], or entropic regularization [32, 34, 31].
@@
Matching procedures based on the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n},\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are invariant to rigid transformations, which is desirable in practice. For instance, in imaging problems, the images of the objects of interest are often obtained from different angles and thus need to be matched up to rigid transformations. However, such procedures are difficult to analyze theoretically and the literature has been mostly studying theoretical guarantees for other matching procedures using $\|X_{i}-Y_{j}\|_{2}$’s, which aim to match the points based on the proximity. [11, 22, 40] establish theoretical results on recovering the ground truth matching for the procedures based on $\|X_{i}-Y_{j}\|_{2}$’s under some noise, and [18, 28, 17] show matching guarantees in the presence of outliers. Although Theorem 2 of [40] concerns a procedure using the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, it requires solving a nontrivial minimization over permutations, which is computationally challenging, and the result does not consider the presence of outliers.
@@
Motivated by the above challenges, this paper theoretically establishes outlier robustness and noise stability of a matching procedure based on the distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, which is easily implementable in practice. The proposed procedure associates each point with the set of distances from it to other points, namely, $X_{i}$ and $Y_{j}$ are associated with the sets of distances to the other points $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$, which we call the distance profiles of $X_{i}$ and $Y_{j}$, respectively. The matching procedure is to find for each $X_{i}$, the point $Y_{j}$ having the most similar distance profile to that of $X_{i}$, where the similarity is measured by a suitable quantity between the sets $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$. The idea of associating a point with suitable invariants based on the distances to other points from it has a long history in the object matching literature, which includes shape contexts [4, 33], integral invariants [25], invariant histograms [8], and the local distribution of distances [26]. However, theoretical results on robustness of such methods are rarely found in the literature. In the context of random graph matching, [12, 13] propose and analyze a degree profile matching procedure based on the adjacency matrices $(A_{i\ell})_{i,\ell=1}^{n}$ and $(B_{j\ell})_{j,\ell=1}^{n}$ of certain correlated random graphs. However, their analysis focuses on random graph models, such as Erdös-Rényi and correlated Wigner models, where the edges are independently generated, namely, $(A_{i\ell})_{i,\ell=1}^{n}$ are mutually independent, which is substantially different from our setting where the distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ are not mutually independent.
@@
We show the outlier robustness and noise stability of practical matching procedures based on distance profiles. The idea of matching points based on invariants like distance profiles has a long history in the literature, proposed and rediscovered multiple times under different names and contexts [4, 33, 25, 8, 26, 12, 13]. However, there has been little understanding of the theoretical properties of such procedures, especially in the presence of outliers and noise. We provide a theoretical analysis showing that under certain probabilistic settings, the proposed matching procedure is successful with high probability even in the presence of outliers and noise. We check the performance of the proposed method using a stylized data example and provide simulation studies to complement the theoretical findings. Lastly, we extend the concept of distance profiles to the abstract setting and connect the proposed matching procedure to the Gromov-Wasserstein distance and its lower bound [26], with a new sample complexity result derived based on the properties of distance profiles. As a result, we contribute to the literature by providing theoretical underpinnings of the matching procedures based on invariants like distance profiles, which have been widely used in practice but have rarely been analyzed theoretically.
@@
For $k\in\mathbb{N}$, let $[k]$ denote $\{1,\ldots,k\}$, $\Delta_{k}=\{(p_{1},\ldots,p_{k})\in\mathbb{R}_{+}^{k}:\sum_{i=1}^{k}p_{i}=1\}$ denote the probability simplex, and $\mathcal{S}_{k}$ denote the set of all permutations on $[k]$. For $a,b\in\mathbb{R}$, let $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. Let $\mathscr{P}(\mathbb{R}^{d})$ denote the set of all Borel probability measures defined on $\mathbb{R}^{d}$ and for any $z\in\mathbb{R}^{d}$, let $\delta_{z}\in\mathscr{P}(\mathbb{R}^{d})$ denote the Dirac measure at $z$. We denote by $W_{p}$ the Wasserstein-$p$ distance for $p\geq 1$. For any measure $\mu$ on a measurable space $\mathcal{X}$ and a measurable function $f\colon\mathcal{X}\to\mathbb{R}$, let $\mu f=\int_{\mathcal{X}}f\,\mathrm{d}\mu$.
@@ 1.1 Distance Profile Matching
We first describe the distance profile matching procedure that we propose to study. As mentioned earlier, we build on the principle of matching points using invariants like distance profiles, which has a long history in the literature [4, 33, 25, 8, 26, 12, 13]. The idea is to associate each point with the set of distances from it to other points, which we call the distance profile, and match the points based on the similarity of the distance profiles. While there are many different procedures incorporating this idea, we are particularly inspired by a simple and easily implementable procedure called degree profile matching [12] that was proposed to match random graphs. We properly modify it to match point clouds, which we summarize in Algorithm 1. We call this procedure the distance profile matching.
@@
First, notice that Algorithm 1 is implementable without knowing the points $X_{i}$’s and $Y_{j}$’s as long as the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are provided. The discrepancy measure $\mathcal{D}$ quantifies the similarity between the distance profiles. Therefore, for each $X_{i}$, the procedure (1.2) finds $Y_{j}$ whose distance profile $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$ is closest to $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$, the distance profile of $X_{i}$, under $\mathcal{D}$.
@@
The output $\pi$ of Algorithm 1 represents the matching: $X_{i}$ is matched to $Y_{\pi(i)}$. Note that $\pi$ may not be injective or surjective, meaning that some points may not be matched or some points may be matched to multiple points. [^1] When $n=m$, Algorithm 1 can be modified to output a permutation (one-to-one correspondence); see Section 3. The motivation behind this is simple: (1) in the presence of outliers, some points can be ignored or matched arbitrarily, and (2) similar points can be matched to the same point. The underlying objects that the points $X_{i}$’s and $Y_{j}$’s approximate, say, probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{d}$, respectively, are not identical in practice. Usually, they coincide up to several outlier parts, which we can formally state as $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$ for some $\lambda\in(0,1)$ and $\rho,\mu_{o},\nu_{o}\in\mathscr{P}(\mathbb{R}^{d})$. Here, $\lambda\rho$ represents the common part of $\mu$ and $\nu$ that we want to match, while $(1-\lambda)\mu_{o}$ and $(1-\lambda)\nu_{o}$ are the outlier parts of $\mu$ and $\nu$, respectively. We are mainly interested in matching $X_{i}$’s and $Y_{j}$’s from the common part, while ignoring or arbitrarily matching the points from the outlier parts.
@@
The threshold $\rho$ is an optional parameter that can be used to obtain the subset $I\subset[n]$ containing the indices $i$’s such that the discrepancy $D(i,\pi(i))$ between the distance profiles of $X_{i}$ and its match $Y_{\pi(i)}$ is below $\rho$. Providing a small threshold leads to a more conservative matching that only focuses on the points that are closely matched, which can be used for other downstream tasks. For instance, one can solve the orthogonal Procrustes problem using the matched pairs $\{(X_{i},Y_{\pi(i)})\}_{i\in I}$ to find the optimal rotation and translation aligning them as closely as possible, which theoretically works with just $|I|=d$ pairs in $\mathbb{R}^{d}$ but practically requires more pairs, say, $|I|=O(d)$, due to potential noise or degeneracy. The obtained transformation can then be used as an initialization of the registration procedures, such as the Iterative Closest Point (ICP) algorithm [5], to find the optimal transformation that aligns the whole point clouds.
@@
The Wasserstein-1 distance between two sets of points in $\mathbb{R}$ admits a closed-form expression based on the corresponding cumulative distribution functions, which can be computed efficiently [30]. In this case, the computational complexity of Algorithm 1 is computed as follows. For Step 1, assuming $n\geq m$, computing the Wasserstein-1 distances between two sets of points in $\mathbb{R}$ for $nm$ pairs takes $O(n^{3}\log(n))$ in total. Meanwhile, Step 2 takes $O(nm\log(m))$ as it requires sorting $m$ numbers $n$ times. Therefore, the overall complexity of Algorithm 1 is $O(n^{3}\log(n))$ assuming $n\geq m$, which is nearly cubic in the number of points $n$; however, as the input dimension of the algorithm is $n^{2}$, we can say that the complexity scales nearly as the input dimension $n^{2}$ raised to the power of $\frac{3}{2}$.
@@
1.2 Stylized Example
Before presenting the main theoretical results, we demonstrate how the proposed distance profile matching procedure performs in practice using a stylized data example. We apply the distance profile matching method to the shape correspondence using CAPOD data [29]. Figure 1(a) and Figure 1(d) show the two 3D images representing dolphins with different poses, from which we obtain point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ as shown in Figure 1(b) and Figure 1(e), respectively, where $n=m=285$ in this example. We want to match the points corresponding to the same part of the dolphins, rather than finding a one-to-one correspondence. As explained earlier, it is reasonable to match similar points to the same point. For instance, there are many points concentrated around the nose—called the rostrum—of the source dolphin, which are expected to be matched to any points concentrated around the rostrum of the target dolphin. Hence, we first divide the dolphins into several components. Here, we use $k$-means clustering with $k=7$, which indeed cluster points based on the body parts. Figure 1(e) shows the target points $Y_{j}$’s colored by $s(j)$’s, where $s(j)\in\{1,\ldots,7\}$ is the index of the component that $Y_{j}$ belongs to. Now, we run Algorithm 1 with $\mathcal{D}=W_{1}$ and obtain the output $\pi\colon[n]\to[m]$. For each $X_{i}$, we verify if the part that it belongs to coincides with the part that its match $Y_{\pi(i)}$ belongs to, which is indexed by $s(\pi(i))$. To this end, Figure 1(b) shows the source points $X_{i}$’s colored by $s(\pi(i))$’s. We can observe that the proposed procedure recovers the corresponding regions mostly correctly. For instance, for $X_{i}$’s corresponding to the rostrum of the source dolphin, their colors $s(\pi(i))$’s (green) indicate that they are matched to the points corresponding to the rostrum of the target dolphin. Similarly, the points corresponding to the head (lime), tail (red), dorsal fin (blue), and peduncle (magenta) are mostly matched correctly, while there are some mismatches, for instance, among the points around the pectoral fin.